Determine the mass of CO2 that contains 7.92 x 1018 O atoms.
Answer #1

ANSWER:

One mole or 44g of CO2 molecules contain 6.022 X 1023 CO2 molecules

1 molecule of CO2 contain 2 atoms of Oxygen

Therefore 44g (1 mole) of CO2 contain 2 X 6.022 X 1023 = 12.044 X 1023 Oxygen atoms

Equating the two

12.044 X 1023 Oxygen atoms = 44g CO2

Or

1 Oxygen atom = 44 / 12.044 X 1023 = 3.65 X 10-23 g CO2

It mean 1 Oxygen atom is present in 3.65 X 10-23 g of CO2

Therefore

7.92 x 1018 atoms of oxygen will be present in 3.65 X 10-23 X 7.92 x 1018 = 28.93 X 10-5 g of CO2